# On Fairness of Low-Rank Adaptation of Large Models

### v1

@@ Abstract
Low-rank adaptation of large models, particularly LoRA, has gained traction due to its computational efficiency. This efficiency, contrasted with the prohibitive costs of full-model fine-tuning, means that practitioners often turn to LoRA and sometimes without a complete understanding of its ramifications. In this study, we focus on fairness and ask whether LoRA has an unexamined impact on utility, calibration, and resistance to membership inference across different subgroups (e.g., genders, races, religions) compared to a full-model fine-tuning baseline. We present extensive experiments across vision and language domains and across classification and generation tasks using ViT-Base, Swin-v2-Large, Llama-2 7B, and Mistral 7B. Intriguingly, experiments suggest that while one can isolate cases where LoRA exacerbates model bias across subgroups, the pattern is inconsistent—in many cases, LoRA has equivalent or even improved fairness compared to the base model or its full fine-tuning baseline. We also examine the complications of evaluating fine-tuning fairness relating to task design and model token bias, calling for more careful fairness evaluations in future work.
@@ 2 Preliminaries & Related Work
Low-Rank Adaptation. LoRA (Hu et al., 2021) is a widely used parameter-efficient fine-tuning algorithm for large models. It proposes to separate out the weight deltas from fine-tuning and approximate them using low-rank matrices; inference then involves forward passing both the (frozen) pre-trained model and the low-rank model deltas, also known as adapters, and summing the activations. Specifically, for a pre-trained weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ with dimensions $d,k$, LoRA approximates its changes from fine-tuning as $\Delta\mathbf{W}\approx\mathbf{B}\mathbf{A}$ where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ with rank $r\ll\min(d,k)$, and thus inference on input $\mathbf{x}\in\mathbb{R}^{d}$ is $\mathbf{W}\mathbf{x}+\mathbf{BA}\mathbf{x}\approx(\mathbf{W}+\Delta\mathbf{W})\mathbf{x}$ if $\Delta\mathbf{W}$ is obtained through full fine-tuning. $\mathbf{A}$ and $\mathbf{B}$ can be updated directly via backpropagation. Typically, implementations of LoRA apply to all query and value matrices of self-attention layers in the pre-trained transformer. To fine-tune for supervised tasks, an additional head is also attached to the last layer of the model.
@@ 3.2 Accuracy
Fairness implications can depend on the quality of pre-trained model. A closer look at Fig. 1 suggests that while LoRA may be considered less fair than full fine-tuning on ViT-Base—by decreased worst subgroup utility on Black group for age classification (1st subplot) and by increased EOD on Asian group for gender classification (3rd subplot)—the tendency disappears when the base model is switched to the more powerful Swin-v2-Large (all else kept the same). This is interesting as it suggests that the fairness properties of LoRA are not only a function of its parameter efficiency and they provide a separation from model pruning where Tran et al. (2022) found that the fairness ramifications persist across model sizes.
@@ 4 Discussions, Limitations, and Future Work
We have presented extensive empirical analyses and found no conclusive evidence that LoRA may exacerbate subgroup fairness compared to full fine-tuning. Does this imply that the parameter efficiency of LoRA is a free lunch? Possibly, but not necessarily—no evidence of unfairness does not imply fairness. While our study aims to be comprehensive, we outline some limitations and directions for future work.
@@
Concluding remarks. Our study sheds light on the fairness properties of low-rank adaptation (LoRA) across architectures, model sizes, datasets, and fairness considerations. In future work, we hope to extend fairness evaluations in generative settings by exploring better experiment design that minimizes the impact of model token bias and/or reliance on reasoning capacity for evaluating generative models. Probing techniques (e.g., Alain & Bengio (2016); Hewitt & Liang (2019); Stoehr et al. (2023); Zou et al. (2023)) emerge as a promising tool to assess models while circumventing their token biases, though the use of additional classifier heads resemble our supervised evaluations. It is also worth exploring and comparing other parameter-efficient methods (e.g., Li & Liang (2021); Liu et al. (2022a)) and their intersection with related techniques such as quantization (Dettmers et al., 2023; Hong et al., 2024) and pruning (Dery et al., 2024; Gromov et al., 2024); this may offer insight whether our findings with LoRA is unique to its algorithmic constructions.
@@ 6 Reproducibility Statement
The code for this project will be open-sourced at https://github.com/kenziyuliu/lora-fairness. All experiments are done using standard machine learning packages, including PyTorch (Paszke et al., 2019), Hugging Face (transformers (Wolf et al., 2019) and PEFT (Mangrulkar et al., 2022)). Distributed training made use of Hugging Face DeepSpeed (Rasley et al., 2020) integration. Other packages used in this work are summarized in the attached requirement.txt file.
@@
7 Acknowledgements
The authors would like to thank Nicolas Papernot for valuable suggestions on an earlier draft of this work. KZL acknowledges support from the Ravi Family Graduate Fellowship from School of Engineering at Stanford University. BI was supported by a Google PhD Fellowship. SK acknowledges support by NSF III 2046795, IIS 1909577, CCF 1934986, NIH 1R01MH116226-01A, NIFA award 2020-67021-32799, the Alfred P. Sloan Foundation, and Google Inc.
@@ A.3 Membership Inference Attacks (MIAs)
The goal of membership inference attacks is to estimate whether a query example is part of a model’s training set as accurately as possible. To this end, consider a dataset space $\mathcal{X}$, label space $\mathcal{Y}$, a real-world distribution $\mathbb{D}$ over $\mathcal{X}\times\mathcal{Y}$, a training dataset $D$ sampled from $\mathbb{D}$, and a training procedure $f_{D}\leftarrow\mathcal{T}(D)$ where $f_{D}$ is a machine learning model that outputs a probability distribution over $\mathcal{Y}$ for any given instance $x\in\mathcal{X}$. The attacker has black-box access to the model $f_{D}$ (known as the target model) and wishes to determine whether the query example $(x,y)\in D$. Following Carlini et al. (2022), evaluating the effectiveness of an MIA can be done by measuring the true positive rate at a low false positive rate. The justification is that being able to confidently infer that just one data point is a member is a much bigger breach in privacy than being $51\%$ confident in the membership of a larger number of datapoints, even though both instances may score the same in other classification metrics such as accuracy or AUROC. In other words, an attack is only an effective privacy breach when it has a low false positive rate.
LiRA. The Likelihood Ratio Attack (LiRA) is an important MIA algorithm due to its effectiveness (Carlini et al., 2022). Here, the attacker trains $N$ shadow models $f_{D_{i}}\leftarrow\mathcal{T}(D_{i})$ where each $D_{i}$ (for $i\in\{1,\dots,N\}$) is randomly sampled from $\mathbb{D}$ (such that for all $i$, the query example $(x,y)\not\in D_{i}$) in order to mimic the behavior of the target model $f_{D}$. For any given model $f:\mathcal{X}\to\mathcal{Y}$, let $f(x)_{y}$ denote the confidence of $f$ on $(x,y)$, in other words, the probability value output by $f(x)$ for label $y$. Although the attacker has no information about $D$, it has complete information about each $D_{i}$ as they were the one who sampled the dataset themselves. Thus, the attacker models a distribution of the confidences $f_{D_{i}}(x)_{y}$ for each shadow model $i$, where $D_{i}$ is sampled from $\mathbb{D}$ and $(x,y)\not\in D_{i}$. From this, the attacker can then compare the target model’s confidence $f_{D}(x)_{y}$ and perform a hypothesis test against the null hypothesis of $(x,y)\not\in D$ (rejecting the null hypothesis and inferring membership whenever the cumulative distribution function of $f_{D}(x)_{y}$ is above some threshold $\tau$). In our implementation, we are performing the “offline” version of the attack, as the traditional “online” approach is infeasibly computationally expensive. For more details about LiRA, we refer the reader to Carlini et al. (2022).
@@ D.5 Effect of Subgroup Size
Figs. A20 and A21 illustrate the effect of increasing group size on utility (e.g., accuracy in the plots) and fairness. We observe that these metrics are not solely dependent on the size of the group:
  • On the D-Lab dataset with language models, accuracy does not consistently increase or decrease with the size of the groups. On the UTK-Face gender dataset with vision models, while there is a general trend of increasing accuracy with larger group sizes, the practical impact of group size on accuracy is limited, since the absolute difference in accuracy across these sizes is marginal.
  • The demographic parity difference and equalized odds difference both exhibit fluctuations across different group sizes without showing a clear trend that correlates with group size. The equalized odds difference tends to decrease as the group size increases in the Llama-2 7B model, but this trend is less clear in the Mistral 7B model (varies in model architectures).
Figure A20: Accuracy and fairness (DPD, EOD) on D-Lab religion subset with subgroups sorted by size.
Figure A21: Accuracy and fairness (DPD, EOD) on UTK-Face gender classification with subgroups sorted by size.

### v2

@@ Abstract
Low-rank adaptation of large models, particularly LoRA, has gained traction due to its computational efficiency. This efficiency, contrasted with the prohibitive costs of full-model fine-tuning, means that practitioners often turn to LoRA without a complete understanding of its ramifications. In this study, we focus on fairness and ask whether LoRA has an unexamined impact on utility, calibration, and resistance to membership inference across different subgroups (e.g., genders, races, religions) compared to a full-model fine-tuning baseline. We present extensive experiments across vision and language domains and across classification and generation tasks using ViT-Base, Swin-v2-Large, Llama-2 7B, and Mistral 7B. Intriguingly, experiments suggest that while one can isolate cases where LoRA exacerbates model bias across subgroups, the pattern is inconsistent—in many cases, LoRA has equivalent or even improved fairness compared to the base model or its full fine-tuning baseline. We also examine the complications of evaluating fine-tuning fairness relating to task design and model token bias, calling for more careful fairness evaluations in future work.
@@ 2 Preliminaries & Related Work
Low-Rank Adaptation. LoRA (Hu et al., 2021) is a widely used parameter-efficient fine-tuning algorithm for large models. It proposes to separate out the weight deltas from fine-tuning and approximate them using low-rank matrices; inference then involves forward passing both the (frozen) pre-trained model and the low-rank model deltas, also known as adapters, and summing the activations. Specifically, for a pre-trained weight matrix $\mathbf{W}\in\mathbb{R}^{d\times k}$ with dimensions $d,k$, LoRA approximates its changes from fine-tuning as $\Delta\mathbf{W}\approx\mathbf{B}\mathbf{A}$ where $\mathbf{B}\in\mathbb{R}^{d\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times k}$ with rank $r\ll\min(d,k)$, and thus inference on input $\mathbf{x}\in\mathbb{R}^{d}$ is $\mathbf{W}\mathbf{x}+\mathbf{BA}\mathbf{x}\approx(\mathbf{W}+\Delta\mathbf{W})\mathbf{x}$ if $\Delta\mathbf{W}$ is obtained through full fine-tuning. $\mathbf{A}$ and $\mathbf{B}$ can be updated directly via backpropagation. Typically, implementations of LoRA apply to all query and value matrices of self-attention layers in a transformer. To fine-tune for supervised tasks, an additional head is also attached to the last layer of the model.
@@ 3.2 Accuracy
Fairness implications can depend on the quality of pre-trained model. A closer look at Fig. 1 suggests that while LoRA may be considered less fair than full fine-tuning on ViT-Base—by decreased worst subgroup utility on Black group for age classification (1st subplot) and by increased EOD on Asian group for gender classification (3rd subplot)—the tendency disappears when the base model is switched to the more powerful Swin-v2-Large (all else kept the same). The results suggest that the fairness properties of LoRA are not only a function of its parameter efficiency; it also implies a separation from model pruning where Tran et al. (2022) found that the fairness ramifications persist across model sizes.
@@ 3.6 Effect of LoRA Rank
3.7 Effect of Subgroup Size
As a control experiment, we investigate the effect of subgroup size on utility and fairness metrics. It is intuitive that subgroups with less samples may be disproportionately affected by the fine-tuning and subsequently experience unfairness. Fig. 10 illustrates the effect of increasing group size on utility (e.g., accuracy in the plots) and fairness. Contrary to the intuition, we did not observe a strong correlation between subgroup size and accuracy and fairness metrics. On the D-Lab dataset for language modeling, accuracy does not consistently increase or decrease with the size of the groups (we defer results on the UTK-Face gender dataset with vision models to Section D.5). The demographic parity difference and equalized odds difference both exhibit fluctuations across different group sizes without showing a clear trend that correlates with group size.
Figure 10: Accuracy and fairness on D-Lab religion subset with subgroups sorted by size.
@@ 4 Discussions, Limitations, and Future Work
We have presented extensive empirical analyses and found no conclusive evidence that LoRA may exacerbate subgroup fairness compared to full fine-tuning. Does this imply that the parameter efficiency of LoRA is a free lunch? Possibly, but not necessarily—no evidence of unfairness does not imply evidence of fairness. While our study aims to be comprehensive, we outline some limitations and directions for future work.
@@
5 Concluding remarks
Our study sheds light on the fairness properties of low-rank adaptation (LoRA) across architectures, model sizes, datasets, and fairness considerations. In future work, we hope to extend fairness evaluations in generative settings by exploring better experiment design that minimizes the impact of model token bias and/or reliance on reasoning capacity for evaluating generative models. Probing techniques (e.g., Alain & Bengio (2016); Hewitt & Liang (2019); Stoehr et al. (2023); Zou et al. (2023)) emerge as a promising tool to assess models while circumventing their token biases, though the use of additional classifier heads resemble our supervised evaluations. It is also worth exploring and comparing other parameter-efficient methods (e.g., ReFT (Wu et al., 2024)), DoRA (Liu et al., 2024)) and their intersection with related techniques such as quantization (Dettmers et al., 2023; Hong et al., 2024) and pruning (Dery et al., 2024; Gromov et al., 2024); this may offer insight whether our findings with LoRA is unique to its algorithmic constructions.
6 Acknowledgements
The authors thank Nicolas Papernot for valuable suggestions on an earlier draft of this work. KZL acknowledges support from the Ravi Family Graduate Fellowship from School of Engineering at Stanford University. BI was supported by a Google PhD Fellowship. SK acknowledges support by NSF IIS 2205329, IIS 2046795, IIS 1909577, CCF 1934986, NIH 1R01MH116226-01A, NIFA award 2020-67021-32799, the Alfred P. Sloan Foundation, Stanford HAI, and Google.
@@ 8 Reproducibility Statement
The code for this project is released at https://github.com/kenziyuliu/lora-fairness. All experiments are done using standard machine learning packages, including PyTorch (Paszke et al., 2019), Hugging Face transformers (Wolf et al., 2019), and Hugging Face PEFT (Mangrulkar et al., 2022). Distributed training made use of Hugging Face DeepSpeed (Rasley et al., 2020) integration. Other packages used in this work are summarized in the attached requirement.txt file.
@@ A.3 Membership Inference Attacks (MIAs)
The goal of membership inference attacks is to estimate whether a query example is part of a model’s training set as accurately as possible. To this end, consider a dataset space $\mathcal{X}$, label space $\mathcal{Y}$, a real-world distribution $\mathbb{D}$ over $\mathcal{X}\times\mathcal{Y}$, a training dataset $D$ sampled from $\mathbb{D}$, and a training procedure $f_{D}\leftarrow\mathcal{T}(D)$ where $f_{D}$ is a machine learning model that outputs a probability distribution over $\mathcal{Y}$ for any given instance $x\in\mathcal{X}$. The attacker wishes to determine, for an instance $x\in\mathcal{X}$ and label $y\in\mathcal{Y}$, whether the query example $(x,y)$ is in $D$. Following Carlini et al. (2022), evaluating the effectiveness of a MIA can be done by measuring the true positive rate at a low false positive rate. The justification is that being able to confidently infer that just one data point is a member is a much bigger breach in privacy than being $51\%$ confident in the membership of a larger number of datapoints, even though both instances may score the same in other classification metrics such as accuracy or AUROC.
LiRA. In the Likelihood Ratio Attack (LiRA) (Carlini et al., 2022), the attacker trains $N$ “shadow models” $f_{D_{i}}\leftarrow\mathcal{T}(D_{i})$ where each $D_{i}$ for $i\in[1,N]$ is randomly sampled from $\mathbb{D}$ such that the query example $(x,y)\not\in D_{i}$ for all $i$. These shadow models are intended to mimic the behavior of the target model $f_{D}$. For any given model $f$, let $f(x)_{y}$ denote the confidence for label $y$ on input $x$ under $f$. Given a fixed example $(x,y)$, the attacker evaluates the model confidence $f_{D_{i}}(x)_{y}$ for all $i$. The attacker then uses the shadow model confidences $f_{D_{i}}(x)_{y}$ to model the distribution of the target model’s confidence $f_{D}(x)_{y}$ under the assumption that $(x,y)\not\in D$ (note that the attacker does not have access to the target dataset $D$, but has full access to each shadow dataset $D_{i}$ as the attacker sampled them). From this, the attacker can then compare the actual value of $f_{D}(x)_{y}$ and perform a hypothesis test against the null hypothesis that $(x,y)\not\in D$ (rejecting the null hypothesis and inferring membership whenever the cumulative distribution function of $f_{D}(x)_{y}$ is above some threshold $\tau$). The version of LiRA described above and implemented in our experiments is the “offline” version of the attack, as the “online” approach is more computationally expensive.
@@ D.5 Effect of Subgroup Size
Fig. A20 illustrates the effect of increasing group size on utility (e.g., accuracy in the plots) and fairness (DPD and EOD). The results support §3.7 that these metrics are not solely dependent on the size of the subgroups.
On the UTK-Face gender dataset with vision models, while there is a general trend of increasing accuracy with larger group sizes, the practical impact of group size on accuracy is limited, since the absolute difference in accuracy across these sizes is marginal.
Figure A20: Accuracy and fairness on UTK-Face gender classification with subgroups sorted by size.